The State of Illinois Cycle Rider Safety Program requires riders to be able to brake from 30 mph (44 ft/s) to 0 in 45 ft. What constant deceleration does it take to do that? 


1
Expert's answer
2021-02-28T07:27:26-0500
"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d},""a=\dfrac{0-(44\ \dfrac{ft}{s})^2}{2\cdot45\ ft}=-21.51\ \dfrac{ft}{s^2}."